Reports 1-1 of 1 Clear search Modify search
MIF (General)
kentaro.komori - 23:41 Tuesday 27 January 2026 (36232) Print this report
DARM sensitivity improvement around 10 Hz

As reported in klog:36219, we have been observing a mysterious improvement in the DARM sensitivity in the 5–20 Hz band immediately after recovering the IFO at room temperature.
This improvement was confirmed by the new calibration applied today (klog:36224).

The attached figure compares the DARM sensitivity achieved at the time of the best sensitivity in O4c with today’s sensitivity obtained with 5-W input power after the calibration work.
An improvement by a factor of 3–10 is observed in the 5–20 Hz range.

Although this significant improvement is clearly real, its origin is not yet understood and should be investigated.
One possible explanation is that the current local damping settings are much better than those in O4c, or that the coupling from other degrees of freedom of some suspensions to DARM is much smaller than it was during O4c.

Images attached to this report
Search Help
×

Warning

×